A mixture of 0.200 moles of N2O and 0.100 moles of O2 is placed in a...

  1. A mixture of 0.200 moles of N2O and 0.100 moles of O2 is placed in a reaction container and allowed to react until equilibrium is established.

2 N2O(g) + O2(g) 4 NO(g)

At equilibrium, 0.196 moles of NO is present. What is the composition of the equilibrium mixture in terms of moles of each substance present?

  1. A 0.178 mole sample of C2H6 and a 0.329 moles sample of H2 are placed in a reaction container and allowed to react until equilibrium is established.

C2H6(g) + H2(g) 2 CH4(g)

At equilibrium, it is found that 0.010 mole of H2 has reacted. What is the composition of the equilibrium mixture in terms of moles of each substance present?
Answer #1

a)

2N2O + O2 <----> 4NO

N2O O2 NO
Initial 0.200 mol 0.100 mol 0 mol
Change -2x -x +4x
Equilibrium (0.200-2x) mol (0.100-x) mol 4x

Now,

4x = 0.196 moles

x = 0.049 mol

At equilibrium,

[N2O] = 0.200 - 2(0.049) = 0.102 mol

[O2] = 0.100 - 0.049 = 0.051 mol

[NO] = 0.196 mol

b)

C2H6 + H2 <------> 2CH4

C2H6 H2 CH4
Initial 0.178 mol 0.329 mol 0 mol
Change -x -x +2x
Equilibrium (0.178-x) mol (0.329-x) mol 2x

Now,

x = 0.010 mol

At equilibrium,

[C2H6] = 0.178 - 0.010 = 0.168 mol

[H2] = 0.329 - 0.010 = 0.319 mol

[CH4] = 2(0.010) = 0.020 mol